# Enhancing Breast Cancer Diagnosis in Mammography: Evaluation and Integration of Convolutional Neural Networks and Explainable AI

### v2

@@ Abstract
The study introduces an integrated framework combining Convolutional Neural Networks (CNNs) and Explainable Artificial Intelligence (XAI) for the enhanced diagnosis of breast cancer using the CBIS-DDSM dataset. Utilizing a fine-tuned ResNet50 architecture, our investigation not only provides effective differentiation of mammographic images into benign and malignant categories but also addresses the opaque ”black-box” nature of deep learning models by employing XAI methodologies, namely Grad-CAM, LIME, and SHAP, to interpret CNN decision-making processes for healthcare professionals. Our methodology encompasses an elaborate data preprocessing pipeline and advanced data augmentation techniques to counteract dataset limitations, and transfer learning using pre-trained networks, such as VGG-16, DenseNet and ResNet was employed. A focal point of our study is the evaluation of XAI’s effectiveness in interpreting model predictions, highlighted by utilising the Hausdorff measure to assess the alignment between AI-generated explanations and expert annotations quantitatively. This approach plays a critical role for XAI in promoting trustworthiness and ethical fairness in AI-assisted diagnostics. The findings from our research illustrate the effective collaboration between CNNs and XAI in advancing diagnostic methods for breast cancer, thereby facilitating a more seamless integration of advanced AI technologies within clinical settings. By enhancing the interpretability of AI-driven decisions, this work lays the groundwork for improved collaboration between AI systems and medical practitioners, ultimately enriching patient care. Furthermore, the implications of our research extend well beyond the current methodologies, advocating for subsequent inquiries into the integration of multimodal data and the refinement of AI explanations to satisfy the needs of clinical practice.
Keywords— Breast Cancer Diagnosis, Mammography, Convolutional Neural Networks, Explainable Artificial Intelligence, ResNet50, Grad-CAM, LIME, SHAP, Data Augmentation, Transfer Learning
@@ 1 Introduction
Breast cancer remains the predominant oncological challenge among women globally, characterized by uncontrolled growth of abnormal cells within the mammary gland. In 2020, this malignancy was diagnosed in approximately 2.3 million women worldwide, culminating in 685,000 fatalities [1]. In response to this global health issue, computer-aided detection (CAD) systems have been developed, leveraging artificial intelligence (AI) algorithms to enhance mammogram interpretation accuracy. These systems facilitate the identification of suspicious regions within digital imaging, subsequently offering diagnostic classifications to assist physicians, thereby acting as an adjunctive opinion [2].
@@
AI methodologies employed in these systems are divided into two principal categories: conventional machine learning algorithms, which use hand-crafted features and deep learning algorithms, which autonomously extract pertinent features from data [3]. Traditional machine learning techniques, such as the K-Nearest Neighbor and Decision Tree algorithms, exhibit limitations in managing voluminous and high-dimensional datasets due to the intricate requirement for robust feature identification and extraction, demanding extensive domain expertise [4]. In contrast, Deep Neural Networks (DNNs), characterized by multiple hidden layers [5], automatically extract features from raw inputs to yield accurate predictions, adjusting to the varied features identified. This automatic feature extraction and classification fusion attribute of deep learning algorithms has been instrumental in attaining unprecedented efficacy in breast cancer detection via imaging modalities [6].
@@
However, the interpretability of DNNs poses a significant challenge due to their ”black-box” nature, an attribute referencing their operational opacity. These networks integrate millions of parameters across numerous layers, thus the complex nature hinders the ability to find an explanation [7]. Although DNNs’ performance in clinical diagnostics is commendable, the lack of transparency and explainability impedes their integration into routine healthcare applications [8]. The emergence of Explainable Artificial Intelligence (XAI) aims to address these issues of transparency and explainability, advocating the creation of AI models that are not only accurate but also comprehensible and trustworthy to users [9].
@@
XAI is characterized as a system that generates models maintaining high accuracy, while simultaneously enabling human users to effectively comprehend and trust these models [10]. Incorporating XAI into a model boosts its credibility, aids in identifying causal relationships within the data, enhances user knowledge, empowers more assured decision-making, and fosters ethical fairness by enhancing transparency [11]. XAI techniques are fundamentally classified into two categories: transparent machine learning and post-hoc models. Transparent models, such as decision trees or K-Nearest Neighbors, are directly interpretable to users due to their inherent simplicity and straightforwardness [12]. Conversely, post-hoc models, encompassing complex deep learning networks like CNNs and RNNs, necessitate additional explanations to comprehend the logic behind their outputs [7],  [13]. Explanations generated by post-hoc methods fall into two primary categories: global and local interpretability. Global interpretability provides a comprehensive view of the model’s functioning, incorporating its overall architecture, training, and data. Local interpretability, however, zeroes in on the rationale for specific predictions, analyzing relevant features and variables [9]. Post-hoc explanatory techniques are also further classified into model-specific and model-agnostic approaches. Model-specific methods, exemplified by Grad-CAM, are devised for particular models and are not transferable across different architectures. Model-agnostic techniques, such as LIME and SHAP, are designed for broader applicability, capable of being deployed across a variety of models. Grad-CAM, LIME, and SHAP are among the leading explainability techniques employed in this research.
@@
The methodology adopted in this research involves a comprehensive preprocessing of mammographic images to enhance their quality for AI analysis, followed by deploying a CNN model for classification. The selection of the CNN architecture is informed by a literature review and the model’s suitability for processing high-dimensional imaging data. To address the challenge of interpretability, this study employs leading XAI techniques, including both model-specific and model-agnostic approaches, to provide insights into the decision-making process of the CNN [9]. The efficacy of these explanatory mechanisms is assessed based on a framework that emphasizes security, transparency, and user comprehensibility [14]. This paper contributes to the growing field of AI in healthcare by providing empirical evidence on the effectiveness of combining CNNs with XAI models for breast cancer detection. It underscores the potential of such integrated approaches to not only enhance diagnostic accuracy but also foster trust and ethical fairness by making AI decisions in healthcare more transparent and understandable [11].
@@ 1.1 Explainability
Interpretability in machine learning (ML) is about comprehending how a model operates, especially when predicting variables. It provides insight into the model’s prediction mechanisms without necessarily unravelling the causal relationships behind those predictions. While interpretability contributes to a model’s overall explainability, not every interpretable model can be deemed fully explainable. It also encompasses the evaluation of hypothetical scenarios, allowing us to speculate on the outcomes had certain inputs been altered. The aim is to achieve a thorough understanding of the ML model, factoring in both visible and hidden elements, to construct a broad explanation of its functionality. The opaque nature of many AI models often obscures the rationale behind their outputs, complicating efforts to understand their decision-making processes. However, interpretability primarily concerns grasping the model’s functional details, whereas explainability extends this by elucidating how the model reacts to new data and the implications of specific changes in its predictions. This distinction becomes particularly significant in healthcare, where the stakes of decision-making are high. A deep comprehension of how AI algorithms derive their suggestions is essential. Without such transparency, healthcare practitioners may hesitate to rely on and embrace AI technologies, fearing the inability to confirm the validity of their recommendations [15]. Thus, explainability and interpretability are essential for verifying the reliability and fairness of AI systems, especially in areas like healthcare, finance, and the legal system where decision-making has significant consequences.
@@
XAI models are typically divided into two main categories: intrinsic models and post-hoc models. Intrinsic models, also known as transparent models, are naturally explainable and understandable due to their simple structure. Common examples of such models include linear regression, logistic regression, and decision trees.
@@
Conversely, post-hoc explainability techniques are further divided into model-specific and model-agnostic methods. Models lacking inherent transparency utilize post-hoc or surrogate models to clarify their decision-making processes. Model-specific explanations are custom-made to fit a model’s unique architecture, with Grad-CAM and saliency maps being notable examples. In contrast, model-agnostic explanations are more versatile and can be applied across different models to shed light on the prediction process, with SHAP and LIME being widely recognized methods in this category. XAI explanations can also be distinguished as either local or global. Local explanations focus on shedding light on individual predictions or outcomes, while global explanations aim to provide an overarching view of how the entire model operates.
The subsequent sections delve into various XAI models, with an emphasis on those most commonly applied alongside ML/Deep Learning (DL) models in the context of breast cancer research. The relationship between ML/DL models and XAI techniques is detailed in Table 1, indicating a prevalent use of SHAP, LIME, and Grad-CAM. Additionally, methods such as layerwise relevance propagation (LRP) and partial dependence plots (PDP) are mentioned for their application in specific studies.
@@
A thorough discussion of XAI models follows, including an overview of relevant research within the breast cancer domain.
@@ 1.2 SHapley Additive exPlanations (SHAP)
Introduced by Lundberg, SHAP [16] offers a coherent framework for the interpretation of model predictions, attributing significance to each feature in the context of a specific prediction. SHAP elucidates the influence of individual features on the output of any machine learning model, providing a tool for both global and local model understanding.
The SHAP methodology is underpinned by game theory, particularly through the lens of Shapley values. It aims to distribute the prediction outcome fairly among the features, akin to dividing a payout among players based on their contribution.
@@
Here, $\phi(f,x)$ symbolizes the SHAP value for a model prediction $f$ with input $x$, where $z^{\prime}$ denotes a subset of features, $M$ the total number of features, and $f_{x}(z^{\prime})$, $f_{x}(z^{\prime}\setminus i)$ represent the model’s predictions for the subset $z^{\prime}$ and for the subset excluding a specific feature, respectively.
@@
    Consistency: The consistency principle states that SHAP values should accurately reflect a feature’s impact. If a feature’s presence or constancy enhances the model’s output, its SHAP value should not decrease. For any two models $f$ and $f^{\prime}$, if
@@
SHAP thus provides a mathematically rigorous framework for fair and consistent feature importance attribution, enhancing the interpretability of machine learning models.
@@ 1.3 Gradient-weighted Class Activation Mapping (Grad-CAM)
Grad-CAM [17], introduced by Selvaraju et al., serves as a visual explanation method for a wide spectrum of Convolutional Neural Network (CNN) models. It operates by identifying and illuminating significant areas within an input image that are pivotal for the model’s class predictions. This is achieved by calculating the gradients of any target class relative to the activations in the final convolutional layer, thereby creating a heatmap of these key regions. Distinguished from CAM, Grad-CAM boasts adaptability across various CNN architectures, enabling model interpretability without necessitating architectural modifications. As a method that discriminates between classes, Grad-CAM conducts gradient calculations and manipulates the feature maps of the concluding convolutional layer to produce meaningful visualizations.
@@
    Gradient Calculation: The commencement of Grad-CAM involves the calculation of the gradient of the loss concerning the final convolutional layer’s activations:
@@
This formula emphasizes the collaborative effect of the weighted activation maps in illuminating the predictive regions of the input image. The ReLU ensures visualization of only those features positively affecting the target class, thereby aiding in the interpretability of the model’s predictive behaviour.
@@ 1.4 Local Interpretable Model-agnostic Explanations (LIME)
Ribeiro et al. introduced Local Interpretable Model-agnostic Explanations (LIME) [18] which generate explanations that are both locally accurate and comprehensible for individual predictions by simulating the model’s decision boundary to be linear in the proximity of the instance being explained. It explicates an instance by fitting an interpretable model to a perturbed dataset around the input instance, shedding light on the prediction mechanisms of the original model.
@@ 2 Related Work
Breast cancer detection often relies on deep learning models to analyze medical imagery, such as mammograms, for diagnosis. These models, though powerful, are frequently viewed as ”black boxes” due to the opaque nature of their decision-making processes. This obscurity can breed mistrust and hinder their adoption in clinical settings. Explainable AI (XAI) seeks to mitigate this problem by making the workings of these models more transparent. For instance, saliency maps are a common XAI technique that illuminates the key features within an image that influence the model’s predictions. This enables healthcare practitioners to better comprehend the basis of a model’s diagnostic suggestions, leading to more informed clinical judgments. Moreover, the methods like SHAP (SHapley Additive exPlanations) further aid in explaining the model predictions for breast cancer detection, offering detailed explanations that healthcare professionals can rely on for making diagnosis decisions. Moncada et al. [19] benchmarked the performance of ML models such as Random Survival Forest (RSF), Survival Support Vector Machine (SSVM), and Extreme Gradient Boosting (XGBoost) against the CPH model for predicting survival outcomes in breast cancer patients. The study focused on female patients from the Netherlands diagnosed with primary invasive non-metastatic breast cancer between 2005 and 2008, highlighting the predictive prowess of XGBoost with a c-index of  0.73, facilitated by SHAP for model explainability.
@@
While in  [20], a hybrid model was introduced that integrates ML with explainable AI (XAI), specifically using XGBoost and SHAP for both prediction and explanation. This approach emphasized the relevance of lifestyle and biological factors, such as high-fat diet and breastfeeding, in breast cancer risk stratification.
@@
Moreover, the integration of radiomics and XAI has opened new avenues in the molecular subtype classification of breast cancer, with studies demonstrating the potential of imaging features to discriminate between luminal and non-luminal subtypes  [21]. SHAP dependence plots have provided insights into the contribution of specific radiomics features to model predictions. It describes how a single feature affects the output of the LASSO_SVM prediction model. The SHAP value is used to estimate the contribution of each feature to the predicted result [22].
@@
For visual explanations of deep learning models’ decisions in medical imaging, Grad-CAM has profound significance in breast cancer diagnosis. By generating heatmaps that highlight key regions in mammograms, Grad-CAM offers clinicians intuitive insights into the features driving model predictions. Its accessibility and ability to bridge the gap between AI-driven diagnostics and clinical practice make Grad-CAM invaluable for enhancing the trust and acceptance of AI systems in healthcare. This visual aspect of Grad-CAM significantly contributes to the interpretability and transparency of medical results, facilitating more informed clinical decisions.
@@
Similarly, in  [25] authors employed DenseNet-169 and EfficientNet-B5 to automate malignancy detection, achieving accuracy levels of 88.1% and 87.9%, respectively. Here, Grad-CAM played a crucial role in highlighting the significant regions leading to malignancy predictions, offering insights into the importance of both the tumour and adjacent areas.
@@
Lou et al.[26] introduced the multi-level global-guided branch-attention network (MGBN) for mass classification in mammograms, incorporating Grad-CAM to verify the network’s accuracy in identifying mass regions. This method not only enhanced the model’s reliability and interpretability but also achieved noteworthy AUC scores of 0.8375 and 0.9311 on the DDSM and INbreast databases, respectively. Building on the theme of enhancing model transparency and efficacy, Wang et al.[27] took a multitask approach with their MIB-Net, which merged classification and tumour segmentation tasks. Utilizing Grad-CAM visualizations, they provided insights into how the model focuses on significant areas across various modalities, showcasing the model’s capability to efficiently prioritize critical information for accurate predictions. Together, these studies highlight the evolving landscape of breast cancer diagnostics, where advanced ML models are increasingly leveraged for their precision, reliability, and interoperability.
@@
LIME provides localized insights into deep learning models’ predictions, including breast cancer diagnosis, by highlighting specific regions of interest in mammograms contributing to the decisions.
@@
The authors in [28] present an automated breast cancer cell image analysis system using the public BreakHis dataset[29]. Utilizing Zernike image moments to extract intricate features from cancer cell images, they employ simple neural networks for binary classification (benign vs. malignant classes). Subsequently, they utilize the LIME technique for explaining the test results, highlighting significant regions responsible for the machine learning algorithm’s decision-making process. This approach enhances the interpretability of test outcomes and justifies the algorithm’s decisions based on input images, addressing the need for explainable AI in histopathology-based breast cancer detection.
@@
Despite the advancements in utilizing Explainable Artificial Intelligence (XAI) techniques for breast cancer detection through medical imaging analysis, a critical gap remains in the quantitative evaluation of these methods. Current literature, including seminal works by Rezazadeh et al.[30], Lamy et al.[31, 32], and others, vividly illustrates the application of XAI to make decision-making processes of deep learning models transparent. However, the absence of a unified framework for the objective measurement and comparison of XAI’s effectiveness limits the ability to ascertain its actual impact on enhancing model interpretability and trustworthiness among healthcare professionals. This gap not only obscures the direct benefits of XAI in clinical settings but also hinders the strategic selection and application of XAI techniques tailored to specific diagnostic challenges in breast cancer detection. It is imperative to establish standardized metrics and benchmarks that enable the rigorous evaluation of XAI techniques across various dimensions, including their influence on diagnostic accuracy, clinician decision-making, and patient outcomes. Such a comprehensive approach would facilitate a deeper understanding of the comparative advantages of different XAI methods, guiding their optimal utilization in clinical practice.

### v3

@@ Abstract
The Deep learning (DL) models for diagnosing breast cancer from mammographic images often operate as ”black boxes,” making it difficult for healthcare professionals to trust and understand their decision-making processes. The study presents an integrated framework combining Convolutional Neural Networks (CNNs) and Explainable Artificial Intelligence (XAI) for the enhanced diagnosis of breast cancer using the CBIS-DDSM dataset. The methodology encompasses an elaborate data preprocessing pipeline and advanced data augmentation techniques to counteract dataset limitations and transfer learning using pre-trained networks such as VGG-16, Inception-V3 and ResNet was employed. A focal point of our study is the evaluation of XAI’s effectiveness in interpreting model predictions, highlighted by utilising the Hausdorff measure to assess the alignment between AI-generated explanations and expert annotations quantitatively. This approach is critical for XAI in promoting trustworthiness and ethical fairness in AI-assisted diagnostics. The findings from our research illustrate the effective collaboration between CNNs and XAI in advancing diagnostic methods for breast cancer, thereby facilitating a more seamless integration of advanced AI technologies within clinical settings. By enhancing the interpretability of AI-driven decisions, this work lays the groundwork for improved collaboration between AI systems and medical practitioners, ultimately enriching patient care. Furthermore, the implications of our research extended well beyond the current methodologies. It encourages further research into how to combine multimodal data and improve AI explanations to meet the needs of clinical practice.
Keywords— Breast Cancer Diagnosis, Mammography, CNNs, XAI, ResNet50, Grad-CAM, LIME, SHAP, Data Augmentation, Transfer Learning
@@ 1 Introduction
Breast cancer remains the predominant oncological challenge among women globally, characterized by the uncontrolled growth of abnormal cells within the mammary gland. In 2020, this malignancy was diagnosed in approximately 2.3 million women worldwide, culminating in 685,000 fatalities [1]. Computer-aided detection (CAD) systems have been developed in response to this global health issue, leveraging artificial intelligence (AI) algorithms to enhance mammogram interpretation accuracy. These systems facilitate the identification of suspicious regions within digital imaging, subsequently offering diagnostic classifications to assist physicians, thereby acting as an adjunctive opinion [2].
@@
AI methodologies employed in these systems are divided into two principal categories: conventional machine learning (ML) algorithms, which use hand-crafted features, and DL, which autonomously extract pertinent features from data [3]. Traditional techniques, such as the K-Nearest Neighbor and Decision Tree algorithms, exhibit limitations in managing voluminous and high-dimensional datasets due to the intricate requirement for robust feature identification and extraction, demanding extensive domain expertise [4]. In contrast, Deep Neural Networks (DNNs), characterized by multiple hidden layers [5], automatically extract features from raw inputs to yield accurate predictions, adjusting to the varied features identified. This automatic feature extraction and classification fusion attribute of DL algorithms has been instrumental in attaining unprecedented efficacy in breast cancer detection via imaging modalities [6].
@@
However, the interpretability of DNNs poses a significant challenge due to their ”black-box” nature, an attribute referencing their operational opacity. These networks integrate millions of parameters across numerous layers; thus, their complex nature hinders the ability to find an explanation [7]. Although DNNs’ performance in clinical diagnostics is commendable, the lack of transparency and explainability impedes their integration into routine healthcare applications [8]. The emergence of XAI aims to address these issues of transparency and explainability, advocating the creation of accurate AI models that are comprehensible and trustworthy to users [9].
@@
XAI is characterized as a system that generates models while maintaining high accuracy and enabling human users to comprehend and trust these models effectively. Incorporating XAI into a model boosts its credibility, aids in identifying causal relationships within the data, enhances user knowledge, empowers more assured decision-making, and fosters ethical fairness by enhancing transparency [10, 11]. XAI techniques are fundamentally classified into two categories: transparent ML and post-hoc models. Transparent models, such as decision trees or K-Nearest Neighbors, are directly interpretable to users due to their inherent simplicity and straightforwardness [12]. Conversely, post-hoc models, encompassing complex DL networks like CNNs and RNNs, necessitate additional explanations to comprehend the logic behind their outputs [7],  [13]. Explanations generated by post-hoc methods fall into two primary categories: global and local interpretability. Global interpretability provides a comprehensive view of the model’s functioning, incorporating its overall architecture, training, and data. Local interpretability, however, zeroes in on the rationale for specific predictions, analyzing relevant features and variables [9]. Post-hoc explanatory techniques are also further classified into model-specific and model-agnostic approaches. Model-specific methods, exemplified by Grad-CAM, are devised for particular models and are not transferable across different architectures. Model-agnostic techniques, such as LIME and SHAP, are designed for broader applicability and are capable of being deployed across a variety of models. Grad-CAM, LIME, and SHAP are among the leading explainability techniques employed in this research.
@@
The methodology adopted in this research involves a comprehensive preprocessing of mammographic images to enhance their quality for AI analysis, followed by deploying a CNN model for classification. The selection of the CNN architecture is informed by a literature review and the model’s suitability for processing high-dimensional imaging data. To address the challenge of interpretability, this study employs leading XAI techniques, including both model-specific and model-agnostic approaches, to provide insights into the decision-making process of CNN [9]. The efficacy of these explanatory mechanisms is assessed based on a framework that emphasizes security, transparency, and user comprehensibility [14]. This paper contributes to the growing field of AI in healthcare by providing empirical evidence on the effectiveness of combining CNNs with XAI models for breast cancer detection. It underscores the potential of such integrated approaches to not only enhance diagnostic accuracy but also foster trust and ethical fairness by making AI decisions in healthcare more transparent and understandable.  [10].
@@ 1.1 Explainability
Interpretability in ML is about comprehending how a model operates, especially when predicting variables. It provides insight into the model’s prediction mechanisms without necessarily unraveling the causal relationships behind those predictions. While interpretability contributes to a model’s overall explainability, not every interpretable model can be deemed fully explainable. It also encompasses the evaluation of hypothetical scenarios, allowing us to speculate on the outcomes had certain inputs been altered. The aim is to achieve a thorough understanding of the ML model, factoring in both visible and hidden elements, to construct a broad explanation of its functionality. The opaque nature of many AI models often obscures the rationale behind their outputs, complicating efforts to understand their decision-making processes. However, interpretability primarily concerns grasping the model’s functional details, whereas explainability extends this by elucidating how the model reacts to new data and the implications of specific changes in its predictions. This distinction becomes particularly significant in healthcare, where the stakes of decision-making are high. A deep comprehension of how AI algorithms derive their suggestions is essential. Without such transparency, healthcare practitioners may hesitate to rely on and embrace AI technologies, fearing the inability to confirm the validity of their recommendations [15]. Thus, explainability and interpretability are essential for verifying the reliability and fairness of AI systems, especially in areas like healthcare, finance, and the legal system where decision-making has significant consequences.
@@
XAI models are primarily divided into two main categories: intrinsic and post-hoc. Intrinsic models, also known as transparent models, are naturally understandable and explainable.
@@
Conversely, post-hoc explainability techniques are further divided into model-specific and model-agnostic methods. Models lacking inherent transparency utilize post-hoc or surrogate models to clarify their decision-making processes. Model-specific explanations are custom-made to fit a model’s unique architecture, with Grad-CAM and saliency maps being notable examples. In contrast, model-agnostic explanations are more versatile and can be applied across different models to explain the prediction process, with SHAP and LIME being widely recognized methods in this category. XAI explanations can also be distinguished as local or global. Local explanations focus on understanding individual decisions made by a model, while global explanations provide insight into the overall behavior and rules of the model across all scenarios.
@@
The subsequent sections delve into various XAI models, with an emphasis on those most commonly applied alongside ML/Deep Learning (DL) models in the context of breast cancer research. A thorough discussion of XAI models follows, including an overview of relevant research within the breast cancer domain.
@@ 1.2 SHapley Additive exPlanations (SHAP)
Introduced by Lundberg, SHAP [16] offers a coherent framework for the interpretation of model predictions, attributing significance to each feature in the context of a specific prediction. SHAP elucidates the influence of individual features on the output of any ML model, providing a tool for both global and local model understanding.
Game theory, particularly as seen through Shapley values, serves as the foundation for the SHAP methodology. It aims to distribute the prediction outcome fairly among the features, akin to dividing a payout among players based on their contribution.
@@
Here, $\phi(f,x)$ symbolizes the SHAP value for a model prediction $f$ with input $x$, where $z^{\prime}$ denotes a subset of features, $M$ the total number of features, and $f_{x}(z^{\prime})$ and $f_{x}(z^{\prime}\setminus i)$ represent the model’s predictions for the subset $z^{\prime}$ and for the subset excluding a specific feature, respectively.
@@
    Consistency: The consistency principle states that SHAP values should accurately reflect a feature’s impact. If a feature’s presence or constancy enhances the model’s output, its SHAP value should not decrease. For any two models, $f$ and $f^{\prime}$, if
@@
SHAP thus provides a mathematically rigorous framework for fair and consistent feature importance attribution, enhancing the interpretability of ML models.
@@ 1.3 Gradient-weighted Class Activation Mapping (Grad-CAM)
Grad-CAM [17], introduced by Selvaraju et al., serves as a visual explanation method for a wide spectrum of CNN models. It operates by identifying and illuminating significant areas within an input image that are pivotal for the model’s class predictions. This is achieved by calculating the gradients of any target class relative to the activations in the final convolutional layer, thereby creating a heatmap of these key regions. Distinguished from CAM, Grad-CAM boasts adaptability across various CNN architectures, enabling model interpretability without necessitating architectural modifications. As a method that discriminates between classes, Grad-CAM conducts gradient calculations and manipulates the feature maps of the concluding convolutional layer to produce meaningful visualizations.
@@
    Gradient Calculation: The first step of Grad-CAM involves the calculation of the gradient of the loss concerning the final convolutional layer’s activations:
@@
This formula emphasizes the collaborative effect of the weighted activation maps on illuminating the predictive regions of the input image. The ReLU ensures visualization of only those features positively affecting the target class, thereby aiding in the interpretability of the model’s predictive behavior.
@@ 1.4 Local Interpretable Model-Agnostic Explanations (LIME)
Ribeiro et al. introduced Local Interpretable Model-Agnostic Explanations (LIME) [18], which generate explanations that are both locally accurate and comprehensible for individual predictions by simulating the model’s decision boundary to be linear in the proximity of the instance being explained. It explicates an instance by fitting an interpretable model to a perturbed dataset around the input instance, shedding light on the prediction mechanisms of the original model.
@@ 2 Related Work
Breast cancer detection often relies on ML/DL models to analyze medical imagery, such as mammograms, for diagnosis. These models, though powerful, are frequently viewed as ”black boxes” due to the opaque nature of their decision-making processes. This obscurity can breed mistrust and hinder their adoption in clinical settings. Explainable AI (XAI) seeks to mitigate this problem by making the workings of these models more transparent. For instance, saliency maps are a common XAI technique that illuminates the key features within an image that influence the model’s predictions. This enables healthcare practitioners to better comprehend the basis of a model’s diagnostic suggestions, leading to more informed clinical judgments. Moreover, methods like SHAP (SHapley Additive exPlanations) further aid in explaining the model predictions for breast cancer detection, offering detailed explanations that healthcare professionals can rely on for making diagnosis decisions. Moncada et al. [19] benchmarked the performance of ML models such as Random Survival Forest (RSF), Survival Support Vector Machine (SSVM), and Extreme Gradient Boosting (XGBoost) against the CPH model for predicting survival outcomes in breast cancer patients. The study focused on female patients from the Netherlands diagnosed with primary invasive non-metastatic breast cancer between 2005 and 2008, highlighting the predictive prowess of XGBoost with a c-index of  0.73, facilitated by SHAP for model explainability.
@@
While in  [20], a hybrid model was introduced that integrates ML with explainable AI (XAI), specifically using XGBoost and SHAP for both prediction and explanation. This approach emphasized the relevance of lifestyle and biological factors, such as a high-fat diet and breastfeeding, in breast cancer risk stratification.
@@
Moreover, the integration of radiomics and XAI has opened new avenues in the molecular subtype classification of breast cancer, with studies demonstrating the potential of imaging features to discriminate between luminal and non-luminal subtypes  [21]. SHAP dependence plots have provided insights into the contribution of specific radiomic features to model predictions. It describes how a single feature affects the output of the LASSO SVM prediction model. The SHAP value is used to estimate the contribution of each feature to the predicted result [22].
@@
For visual explanations of DL models’ decisions in medical imaging, Grad-CAM has profound significance in breast cancer diagnosis. By generating heatmaps that highlight key regions in mammograms, Grad-CAM offers clinicians intuitive insights into the features driving model predictions. Its accessibility and ability to bridge the gap between AI-driven diagnostics and clinical practice make Grad-CAM invaluable for enhancing the trust and acceptance of AI systems in healthcare. This visual aspect of Grad-CAM significantly contributes to the interpretability and transparency of medical results, facilitating more informed clinical decisions.
@@
Similarly, in  [25] authors employed DenseNet-169 and EfficientNet-B5 to automate malignancy detection, achieving accuracy levels of 88.1% and 87.9%, respectively. Here, Grad-CAM played a crucial role in highlighting the significant regions leading to malignancy predictions, offering insights into the importance of both the tumor and adjacent areas.
@@
Lou et al.[26] introduced the multi-level global-guided branch-attention network (MGBN) for mass classification in mammograms, incorporating Grad-CAM to verify the network’s accuracy in identifying mass regions. This method not only enhanced the model’s reliability and interpretability but also achieved noteworthy AUC scores of 0.8375 and 0.9311 on the DDSM and INbreast databases, respectively. Building on the theme of enhancing model transparency and efficacy, Wang et al.[27] took a multitask approach with their MIB-Net, which merged classification and tumor segmentation tasks. Utilizing Grad-CAM visualizations, they provided insights into how the model focuses on significant areas across various modalities, showcasing the model’s capability to efficiently prioritize critical information for accurate predictions. Together, these studies highlight the evolving landscape of breast cancer diagnostics, where advanced ML models are increasingly leveraged for their precision, reliability, and interoperability.
@@
LIME provides localized insights into the models’ predictions, including breast cancer diagnosis, by highlighting specific regions of interest in mammograms contributing to the decisions.
@@
The authors in [28] present an automated breast cancer cell image analysis system using the public BreakHis dataset[29]. Utilizing Zernike image moments to extract intricate features from cancer cell images, they employ simple neural networks for binary classification (benign vs. malignant classes). Subsequently, they utilize the LIME technique for explaining the test results, highlighting significant regions responsible for the ML algorithm’s decision-making process. This approach enhances the interpretability of test outcomes and justifies the algorithm’s decisions based on input images, addressing the need for explainable AI in histopathology-based breast cancer detection.
@@
Despite the advancements in utilizing XAI techniques for breast cancer detection through medical imaging analysis, a critical gap remains in the quantitative evaluation of these methods. Current literature, including seminal works by Rezazadeh et al.[30], Lamy et al.[31, 32], and others, vividly illustrates the application of XAI to make the decision-making processes of DL models transparent. However, the absence of a unified framework for the objective measurement and comparison of XAI’s effectiveness limits the ability to ascertain its actual impact on enhancing model interpretability and trustworthiness among healthcare professionals. This gap not only obscures the direct benefits of XAI in clinical settings but also hinders the strategic selection and application of XAI techniques tailored to specific diagnostic challenges in breast cancer detection. It is imperative to establish standardized metrics and benchmarks that enable the rigorous evaluation of XAI techniques across various dimensions, including their influence on diagnostic accuracy, clinician decision-making, and patient outcomes. Such a comprehensive approach would facilitate a deeper understanding of the comparative advantages of different XAI methods, guiding their optimal utilization in clinical practice.